### v3

# Yunrui Li,a Hao Xu,a and Pengyu Honga

[…]
[Uncaptioned image] AI-enabled prediction of NMR spectroscopy: Deducing 2-D NMR of carbohydrate
Yunrui Li,a Hao Xu,a and Pengyu Honga
[Uncaptioned image] In the dynamic field of nuclear magnetic resonance (NMR) spectroscopy, artificial intelligence (AI) has ushered in a transformative era for molecular studies. AI-driven NMR prediction, powered by advanced machine learning and predictive algorithms, has fundamentally reshaped the interpretation of NMR spectra. This innovation empowers us to forecast spectral patterns swiftly and accurately across a broad spectrum of molecular structures. Furthermore, the advent of generative modeling offers a groundbreaking approach, making it feasible to make informed prediction of 2D NMR from chemical language (such as SMILES, IUPAC Name). Our method mirrors the multifaceted nature of NMR imaging experiments, producing 2D NMRs for the same molecule based on different conditions, such as solvents and temperatures. Our methodology is versatile, catering to both monosaccharide-derived small molecules, oligosaccharides and large polysaccharides. A deeper exploration of the discrepancies in these predictions can provide insights into the influence of elements such as functional groups, repeating units, and the modification of the monomers on the outcomes. Given the complex nature involved in the generation of 2D NMRs, our objective is to fully leverage the potential of AI to enhance the precision, efficiency, and comprehensibility of NMR spectral analysis, ultimately advancing both the field of NMR spectroscopy and the broader realm of molecular research.
footnotetext: a Department of Computer Science, Brandeis University, USA; E-mail: hongpeng@brandeis.edufootnotetext: ‡ These authors contributed equally to this work.
1 Introduction
Nuclear magnetic resonance (NMR) spectroscopy has emerged as a versatile tool with widespread applications across diverse scientific domains, including chemistry, environmental science, food science, material science, and drug discovery by unraveling molecular dynamics and structures.1, 2, 3 The primary information of an NMR spectrum arises from the chemical shift, which is determined by the local environment of a nucleus and influenced by interactions through chemical bonds and space. This mechanism yields unique “fingerprints” corresponding to diverse functional groups or molecular motifs, thereby facilitating the streamlined deduction of atomic connectivity and arrangement.
Interpreting NMR spectra requires following essential guidelines, often referred to as “rules of thumb”, where specific chemical shifts are associated with distinctive functional groups.4 The determination of molecular structures from varying chemical shifts on NMR spectra generally requires the expertise of experienced organic chemists. To facilitate the interpretation of NMR spectra, significant efforts have been directed towards computational simulation of NMR spectra.5 Early computational approaches, like the Hierarchically Ordered Spherical Environment (HOSE) codes 6, aim to encapsulate atom neighborhoods in concentric spheres, utilizing a nearest-neighbor approach to predict NMR shift values. A recent HOSE approach 7 yields mean absolute errors (MAEs) of 3.52 ppm for 13C NMR and 0.29 ppm for 1H NMR on the nmrshiftdb2 8 dataset. Concurrently, significant efforts have been devoted to the ab initio calculation of NMR properties.9, 10 Density Functional Theory (DFT)-based methods were developed for certain small organic molecules, achieving MAEs of 2.9 ppm for 13C NMR and 0.23 ppm for 1H NMR.11 However, the accuracy of these DFT-based methods relies heavily on the choice of the basis functions, which often require meticulous case-by-case manual tuning for each molecule. Moreover, the time-intensive nature of DFT calculations limits their applications to comprehensive and large datasets. Recently, the rise of Graph Neural Networks (GNN) and their successes in predicting molecular properties 12, 13, 14, 15, 16 has prompted initiatives to employ GNNs for predicting peaks in NMR spectra. The application of GNN to molecules is intuitive, as a molecular structure can be naturally represented as a graph, with each atom as a node and its chemical bonds as edges. On 1D NMR data, a GNN-based model achieves MAEs of 1.355 ppm for 13C NMR and 0.224 ppm for 1H NMR on the nmrshiftdb2 dataset.17 While considerable efforts have been made in developing predictive models for 1D NMR, the prediction of 2D NMR remains underexplored.
Heteronuclear Single Quantum Coherence (HSQC) spectroscopy 18, a sophisticated 2D NMR technique, serves as a pivotal tool for elucidating atomic connectivity within complex molecules where conventional 1D NMR may prove insufficient.19, 20 By correlating the chemical shifts of hydrogen nuclei with those of heteronuclear nuclei, typically carbon or nitrogen, via scalar coupling interactions, HSQC facilitates the comprehensive mapping of interatomic connections within a molecule. This mapping yields crucial insights into chemical bonding, molecular conformation, and intramolecular interactions. A notable stride in this domain utilizes the ML approach to establish correlations between DFT-simulated HSQC spectra and empirical data to identify molecules.21 However, the accurate prediction of HSQC spectra using ML techniques remains elusive. These obstacles primarily arise from the following factors: the scarcity of annotated datasets for training ML models, difficulties in handling the inherent sparsity of HSQC spectra which complicates feature representation, the computational demands required for accurate full-spectrum prediction, and the need for an in-depth comprehension of molecular structures.
In light of the aforementioned challenges and opportunities in interpreting HSQC spectra, we propose a novel solvent-aware machine learning model designed to predict HSQC cross peaks based on molecular structures, with the capability of peak assignment for experimental spectra, illustrated in Figure 1. Alongside a Graph Neural Network (GNN) module capturing structural nuances, the model incorporates a solvent encoder to effectively account for the impact of solvent environments on chemical shifts, which is essential for delivering accurate cross peak prediction and peak assignment of HSQC spectra. To tackle the lack of annotated HSQC data, we designed a two-step training process. First, the model is pre-trained on a labeled 1D NMR dataset, enabling it to learn a wide range of C–H interactions. Then, we implement an unsupervised strategy that uses the unlabeled HSQC dataset to refine the model’s ability to accurately discern and label HSQC cross peaks. The model is thoroughly evaluated and compared to traditional tools like ChemDraw22 and Mestrenova23 using expert-annotated test datasets. It demonstrates superior accuracy across all molecular weight categories compared to these traditional tools.
Fig.  1: Illustration of HSQC prediction and alignment pipeline. The model takes a molecular structure and derives its atomic representations. The final output of the model has both the HSQC cross peaks and atom alignment.
2 Methods
Our model integrates several components to facilitate the understanding of NMR data through innovative computational approaches. At its core, the model employs a Graph Neural Network (GNN) for encoding molecular features alongside a solvent encoder for capturing the effects of different solvents on NMR chemical shifts. The combination of these components allows the model to generate detailed predictions about molecular behavior in various environments, crucial for interpreting complex NMR data.
The GNN, as illustrated in Figure 2 is designed to learn detailed embeddings of atoms within a molecule, capturing both local and broader chemical environments which are critical for predicting NMR chemical shifts. This is achieved by representing molecules as graphs where nodes are atoms and edges represent bonds, with various attributes such as atomic type and bond characteristics enhancing the model’s understanding of molecular structure.
The model’s ability to integrate detailed molecular structure with solvent effects makes it a powerful tool for analyzing NMR data. This holistic approach is crucial for advancing our understanding of molecular interactions in different chemical environments, which is key for numerous applications in chemistry and related fields.
Fig.  2: Illustration of message passing and node representation updates in a GNN layer. (A) For a given center atom in a molecular graph, the local environment contains the neighborhood atom that is directly bonded to the center atom. (B) For a center node ($v$), initial random representations are assigned to this node, its neighboring nodes ($u_{1}$, $u_{2}$, $u_{3}$), and their connecting bonds ($e_{u_{1},v}$, $e_{u_{2},v}$, $e_{u_{3},v}$). (C) Message Passing and node update. Representations of all neighboring nodes and edges are aggregated and integrated to form a message to the center node. The representation of the center node is then updated to incorporate this message and information from its previous state.
3 Results and discussion
3.1 Data
Two datasets are used in the training process. The first one is the NMRShiftDB28 dataset which contains $\sim$24,000 annotated NMR spectra collected from 22,663 distinct molecules. The second dataset is collected from HMDB 24 and CH-NMR-NP 25, divided into a training dataset containing $\sim$19,000 experimental HSQC spectra and a validation dataset containing $\sim$5,000 HSQC spectra.
3.2 Pre-training performance
For 1D NMR data, our model simultaneously predict ${}^{1}\text{H}$ and ${}^{13}\text{C}$ NMR spectra using a single model framework. Our training and validation data includes shifts for ${}^{13}\text{C}$, ${}^{1}\text{H}$, or both, with the majority of the samples containing only ${}^{13}\text{C}$, a smaller subset of samples with both ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts and only a few samples containing ${}^{1}\text{H}$ shifts. Despite the small portion of the ${}^{1}\text{H}$ shifts in the dataset, our model achieved impressive mean absolute errors (MAEs) of 0.210 ppm for ${}^{1}\text{H}$ NMR prediction and 2.228 ppm for ${}^{13}\text{C}$ NMR prediction, demonstrating the effectiveness of our approach. The success of our model can be attributed to GNN and message passing, which allows the model to learn ${}^{1}\text{H}$ atom features not only from directly prediction the ${}^{1}\text{H}$ shifts, but also from the connecting ${}^{13}\text{C}$ features through message-passing mechanisms. This approach helps the model to implicitly learn the C–H interactions which laid a solid foundation for the HSQC training in the next step. Using this method, our framework surpasses the data limitation from experiments where ${}^{13}\text{C}$ signals are easier to obtain than ${}^{1}\text{H}$, or vice versa. It offers a flexible procedure to utilize experimental data and learn molecular dynamics from them as much as possible. Overall, these results highlight the potential of multitasking training in enhancing the accuracy and performance of NMR prediction models.
Our 1D NMR data training incorporates two specifically designed features to optimize the model for 2D NMR predictions, albeit without immediate benefits for 1D NMR outcomes. Firstly, in our dataset for 1D NMR, solvent labels are scarcely provided. Consequently, we handle the solvent class as “unknown” during training. This approach yields no observable enhancement in performance compared to a model that does not account for solvent information. However, it facilitates a seamless adaptation to the HSQC dataset, characterized by its varied solvent labels, preparing the model for a more nuanced solvent-aware prediction in 2D NMR scenarios. Additionally, due to the differing distribution of the HSQC dataset (76.34% small molecules and 90.33% non-saccharides) compared to the 1D NMR datasets (98.80% small molecules and 99.95% non-saccharides), we anticipate challenges in transferring learning or fine-tuning the model for HSQC prediction. To address this, we intend to use fewer epochs, specifically 200 epochs, as opposed to the 500 epochs used in MPNN work 17. To reduce model overfitting on 1D data, we also left 20% data as validation and only 80% for training, compared to 5% and 95% data split in the MPNN work. This modification, while possibly resulting in slightly elevated MAEs for 1D predictions, enhances the model’s flexibility for HSQC fine-tuning and aligns with our strategic objective of creating a solvent-aware model suited for HSQC spectrum analysis. This balance between predictive accuracy and computational expediency is pivotal, enhancing our model’s overall feasibility and effectiveness for intricate HSQC predictions. Table 1 shows the model’s performance compared to benchmarak methods on 1D NMR dataset.
Table 1: MAEs for representative methods of 1D NMR prediction
Method 1H MAEs (ppm) 13C MAEs(ppm)
HOSE 8 0.29 3.52
DFT 11 0.23 2.9
MPNN 17 0.224 1.355
GNN + solvent 0.210 2.228
3.3 Convergence during self-training
Self-training is an iterative process aimed at minimizing validation errors while avoiding overfitting. We conducted a few iterations, as illustrated in Figure 3. Across five iterations, we observed a significant decrease in validation errors, particularly in the initial stages. In the first iteration, MAEs of C-shifts decreased by approximately 1.1 ppm, and MAEs of H-shifts decreased by about 0.12 ppm. Subsequently, in the second iteration, there was a decrease in MAEs of C-shifts by roughly 0.3 ppm and MAEs of H-shifts by around 0.1 ppm. After 5 iterations, the improvement becomes marginal between iterations. Consequently, we conclude that the process has converged. The MAE of model prediction error is shown in Table 2. Upon the convergence, the model is capable of producing accurate HSQC predictions while concurrently generating the atom alignment for each C–H bond. In Figure 4, we demonstrate an example of HSQC prediction and peak assignments.
Fig.  3: Error curve for ${}^{13}\text{C}$ and ${}^{1}\text{H}$ during iterative self-training phases. The pre-trained error shows direct application of the pre-trained model on 1D data to predict HSQC shifts. After 4 rounds of iterative self-training, the error curve gradually flattens, indicating the model has reaches satisfactory performance.
Table 2: MAEs of 2D NMR prediction after convergence
Dimension Training MAEs (ppm) Validation MAEs (ppm)
C-shifts 1.638 1.997
H-shifts 0.131 0.161
Fig.  4: Example of model usage on HSQC prediction and peak assignment. The molecule is shown in the left, where each C–H bond is labelled. The model’s HSQC spectrum prediction is shown in the right, with the model assigning accurate labels to these predictions. Notably, the model adeptly recognizes and accounts for symmetry within the molecule, assigning identical C–H shifts to symmetrical bond sets. This exemplifies the model’s advanced capability to discern and reflect molecular conformation in its predictions accurately.
3.4 Post-analysis of model performance
In this section, we highlight the model’s effectiveness across different molecular categories, emphasizing its performance on medium to large molecules and saccharides.
3.4.1 Performance regarding molecular weight
The interpretation of HSQC spectra for medium to large molecules often presents significant challenges, as their considerable size introduces greater structural complexity and various isomeric configurations. This complexity introduce more intramolecular interactions, and often results in crowded NMR spectra with overlapping signals, making it difficult to resolve individual peaks. Additionally, the increased number of spin-spin interactions within large molecules demands more advanced NMR techniques to achieve sufficient resolution. The longer relaxation times associated with larger molecules slow down experimental procedures, requiring more sensitive NMR equipment. Furthermore, solubility issues can lead to weak signals, complicating spectral analysis even further. Therefore, there is a pressing need for a model proficient in predicting and analyzing HSQC spectra for medium to large molecules.
As demonstrated in Figure 5, our model achieves significant accuracy in predicting shifts for both medium and large molecules. We categorize the dataset into three segments based on molecular weight (MW), utilizing 500 and 1000 daltons as the dividing thresholds. For small molecules (MW < 500 daltons), the HSQC prediction yields a MAE of 1.75 for ${}^{13}\text{C}$ shifts and 0.14 for ${}^{1}\text{H}$ shifts. For medium molecules (500 < MW < 1000 daltons), the HSQC prediction results in an MAE of 1.81 for ${}^{13}\text{C}$ shifts and 0.16 for ${}^{1}\text{H}$ shifts. For large molecules (MW > 1000 daltons), the HSQC prediction shows an MAE of 2.18 for ${}^{13}\text{C}$ shifts and 0.18 for ${}^{1}\text{H}$ shifts. The predictions for small and medium molecules exhibit comparable accuracy, with a slight reduction in precision for larger molecules, likely due to their rarity in the training dataset. Notably, large molecules constitute merely 2% of our dataset, reflecting the scarcity and difficulty in obtaining high-quality 2D NMR data for such molecules. Overall, our model demonstrates stable performance across different sizes of molecules, underscoring the effectiveness and promise of our framework in delivering precise and reliable 2D NMR shift predictions.
Fig.  5: Model performance comparison between small, medium amd large molecules by molecular weight. The model performs equally well between small and medium molecules, with a marginally reduced precision for larger molecules.
Fig.  6: Model performance comparison between saccharides and non-saccharides. The model performs equally well in both groups.
3.4.2 Performance on saccharides
Fig.  7: Examples of model performance on saccarides. The orange dots indicate the HSQC prediction by our model and the blue triangles indicates the observed signals from raw experiment data. For each example, the Mean Absolute Error (MAE) is shown in the lower right of the plot. Our model’s prediction is able to account for every signal that is observed from the experiments.
Saccharides, or carbohydrates, play critical roles in biological processes, serving as the primary source of energy for living organisms and functioning as essential structural components in cells and tissues.26, 27 Their ubiquitous presence and diverse functions underscore their importance in maintaining the delicate balance of life. From the glucose molecules that fuel cellular respiration to the intricate glycoproteins that mediate cellular recognition and signaling, carbohydrates form an indispensable cornerstone of biological systems.28, 29 Moreover, their dynamic interplay with proteins, lipids, and nucleic acids orchestrates a myriad of physiological processes, ranging from cell-cell communication to immune response modulation. 30 In essence, saccharides actively contribute to the complexity of life through their multifunctional roles and interactions within biological systems.
However, elucidating the structures of saccharides presents a formidable challenge due to their inherent structural complexity and diversity. This complexity arises from the diverse arrangements of monosaccharide units, varied anomeric configurations, and variable glycosidic linkages. Additionally, carbohydrates often lack the requisite crystallinity, unlike the well-defined crystalline structures of small molecules or proteins, which hinders the acquisition of high-resolution X-ray diffraction data. Consequently, NMR spectroscopy, particularly through techniques such as HSQC, has emerged as an indispensable tool in unraveling the detailed structures of carbohydrates.31
Our model demonstrates excellent performance in predicting HSQC spectra for saccharides, achieving a level of accuracy comparable to that observed for non-saccharide molecules, as illustrated in Figure 6. For non-saccharides, the HSQC prediction yields a MAE of 1.765 for ${}^{13}\text{C}$ shifts and 0.138 for ${}^{1}\text{H}$ shifts. For saccharides, the HSQC prediction yields a MAE of 1.808 for ${}^{13}\text{C}$ shifts and 0.163 for ${}^{1}\text{H}$ shifts. Furthermore, some examples are presented in Figure 7. These saccharide molecules feature multiple ring structures and numerous stereogenic centers, contributing to the intricate nature of their HSQC spectra. Despite these inherent challenges, our proposed model exhibits exceptional accuracy in predicting the spectra for these molecules. This robust performance underscores the model’s capacity to navigate the complexities associated with saccharide molecules, thereby emphasizing its versatility and effectiveness across various applications in the field.
3.5 Comparison to traditional tools
In organic chemistry, the simulation of HSQC spectra holds significant importance for analyzing experimental HSQC sepctra. It enables researchers to assign crosspeaks to C–H bonds in the target molecule. Traditional approaches, encompassing web-based tools like nmrdb.org 32, software solutions such as ChemDraw22, and Mestronova23, have long served as go-to resources for this task. In our investigation, we present two compounds (One is simple, the other is more complex.) for the comparison of classical methods with our innovative model.
Our model demonstrates clear advantages over traditional tools, as shown in Figure 8. For the simple molecule (a), traditional methods like ChemDraw, Mestronova, and nmrdb.org exhibit moderate performance, with MAEs for 13C shifts ranging from 1.35 to 3.47 ppm, and for 1H shifts ranging from 0.18 to 0.27 ppm. In contrast, our model’s predictions closely align with experimental observations, with MAEs of only 0.35 ppm for ${}^{13}\text{C}$ and 0.07 ppm for ${}^{1}\text{H}$. This disparity is even more pronounced for the larger molecule (b), where traditional methods show significantly larger MAEs ranging from 9.63 to 20.02 ppm for 13C shifts and from 0.74 to 1.19 ppm for 1H shifts. In contrast, our model achieves MAEs of only 0.67 ppm for ${}^{13}\text{C}$ shifts and 0.07 ppm for ${}^{1}\text{H}$ shifts.
Fig.  8: Performance comparison between proposed GNN model and traditional method. For each example (a) and (b), the observed signals from experiments are annotated by blue dots and the prediction by each method is highlighted in orange. The prediction error, measured by MAE, is shown in the bottom right on each plot. For a small molecule (a), all methods perform within reason but our model offers the most precise performance. For a large molecule (b), the advantage of our model becomes more evident, as the traditional methods become less accurate in analyzing complicated conformations.
4 Conclusions and Discussions
In this work, we introduced a novel approach for predicting HSQC spectra, aimed at addressing the challenges arising from significant computation cost, limited data availability and labor-intensive annotation. Our framework integrates a GNN and solvent impact to address these challenges effectively. The GNN facilitates the generation of atomic embeddings that encapsulate both local and global chemical environments surrounding each atom. These atomic embeddings capture intricate chemical properties and interactions that profoundly influence chemical shifts. Subsequently, multilayer perceptron (MLP) layers are employed to predict C–H shifts based on these refined atom embeddings. This approach effectively captures solvent effects, thereby enabling accurate predictions of chemical shifts.
For future directions, we aim to broaden the application of GNN-based NMR predictions to complex biological molecules, including proteins, nucleic acids, and large polysaccharides. Accurate prediction of NMR spectra for these molecules holds the potential to deepen our understanding of their structure-function relationships, dynamics, and interactions with other molecules. This advancement could significantly contribute to accelerating drug discovery and biochemical research efforts.
Data Availability
The codes will be made publicly available upon the date of publication, and the datasets will be available upon request.
Conflicts of interest
There are no conflicts of interest to declare.
Acknowledgements
This work was supported by GlycoMIP, a National Science Foundation (NSF) Materials Innovation Platform funded through Cooperative Agreement DMR-1933525, as well as NSF OAC 1920147.
Notes and references
  • Gunther and Gunther 1994 H. Gunther and H. Gunther, NMR spectroscopy: basic principles, concepts, and applications in chemistry, John Wiley & Sons Chichester, UK, 1994.
  • Claridge 2016 T. D. Claridge, High-resolution NMR techniques in organic chemistry, Elsevier, 2016, vol. 27.
  • Yu et al. 2021 H.-Y. Yu, S. Myoung and S. Ahn, Magnetochemistry, 2021, 7, 121.
  • Jonas and Kuhn 2019 E. Jonas and S. Kuhn, Journal of cheminformatics, 2019, 11, 1–7.
  • Jonas et al. 2022 E. Jonas, S. Kuhn and N. Schlörer, Magnetic Resonance in Chemistry, 2022, 60, 1021–1031.
  • Bremser 1978 W. Bremser, Analytica Chimica Acta, 1978, 103, 355–365.
  • Kuhn and Johnson 2019 S. Kuhn and S. R. Johnson, ACS omega, 2019, 4, 7323–7329.
  • Steinbeck et al. 2003 C. Steinbeck, S. Krause and S. Kuhn, Journal of chemical information and computer sciences, 2003, 43, 1733–1739.
  • Lodewyk et al. 2012 M. W. Lodewyk, M. R. Siebert and D. J. Tantillo, Chemical Reviews, 2012, 112, 1839–1862.
  • Willoughby et al. 2014 P. H. Willoughby, M. J. Jansma and T. R. Hoye, Nature protocols, 2014, 9, 643–660.
  • Wiitala et al. 2006 K. W. Wiitala, T. R. Hoye and C. J. Cramer, Journal of Chemical Theory and Computation, 2006, 2, 1085–1092.
  • Wieder et al. 2020 O. Wieder, S. Kohlbacher, M. Kuenemann, A. Garon, P. Ducrot, T. Seidel and T. Langer, Drug Discovery Today: Technologies, 2020, 37, 1–12.
  • Zhang et al. 2022 Z. Zhang, L. Chen, F. Zhong, D. Wang, J. Jiang, S. Zhang, H. Jiang, M. Zheng and X. Li, Current Opinion in Structural Biology, 2022, 73, 102327.
  • Fang et al. 2022 X. Fang, L. Liu, J. Lei, D. He, S. Zhang, J. Zhou, F. Wang, H. Wu and H. Wang, Nature Machine Intelligence, 2022, 4, 127–134.
  • Wang et al. 2023 Y. Wang, S. Chen, G. Chen, E. Shurberg, H. Liu and P. Hong, Informatics, 2023, p. 8.
  • Xu et al. 2024 H. Xu, Z. Zhou and P. Hong, arXiv preprint arXiv:2401.17615, 2024.
  • Kwon et al. 2020 Y. Kwon, D. Lee, Y.-S. Choi, M. Kang and S. Kang, Journal of chemical information and modeling, 2020, 60, 2024–2030.
  • Bodenhausen and Ruben 1980 G. Bodenhausen and D. J. Ruben, Chemical Physics Letters, 1980, 69, 185–189.
  • Bross-Walch et al. 2005 N. Bross-Walch, T. Kühn, D. Moskau and O. Zerbe, Chemistry & biodiversity, 2005, 2, 147–177.
  • Li and Kang 2020 Q. Li and C. Kang, Molecules, 2020, 25, 2974.
  • Zanardi and Sarotti 2015 M. M. Zanardi and A. M. Sarotti, The Journal of organic chemistry, 2015, 80, 9371–9378.
  • Mills 2006 N. Mills, ChemDraw Ultra 10.0 CambridgeSoft, 100 CambridgePark Drive, Cambridge, MA 02140. www. cambridgesoft. com. Commercial Price: 1910fordownload, 2150 for CD-ROM; Academic Price: 710fordownload, 800 for CD-ROM., 2006.
  • Willcott 2009 M. R. Willcott, MestRe nova, 2009.
  • Wishart et al. 2022 D. S. Wishart, A. Guo, E. Oler, F. Wang, A. Anjum, H. Peters, R. Dizon, Z. Sayeeda, S. Tian, B. L. Lee et al., Nucleic acids research, 2022, 50, D622–D631.
  • 25 J-resonance, NMR Database, https://www.j-resonance.com/en/nmrdb/.
  • Guillén et al. 2010 D. Guillén, S. Sánchez and R. Rodríguez-Sanoja, Applied microbiology and biotechnology, 2010, 85, 1241–1249.
  • Yu et al. 2018 Y. Yu, M. Shen, Q. Song and J. Xie, Carbohydrate polymers, 2018, 183, 91–101.
  • Dashty 2013 M. Dashty, Clinical biochemistry, 2013, 46, 1339–1352.
  • Fadaka et al. 2017 A. Fadaka, B. Ajiboye, O. Ojo, O. Adewale, I. Olayide and R. Emuowhochere, Journal of Oncological Sciences, 2017, 3, 45–51.
  • Sun et al. 2016 L. Sun, D. R. Middleton, P. L. Wantuch, A. Ozdilek and F. Y. Avci, Glycobiology, 2016, 26, 1029–1040.
  • Fels and Bunzel 2022 L. Fels and M. Bunzel, Magnetic Resonance in Chemistry, 2022, 60, 692–701.
  • Banfi and Patiny 2008 D. Banfi and L. Patiny, Chimia, 2008, 62, 280–280.